OpenStudy (anonymous):

A model rocket is launched with an initial upward velocity of201ft/s . The rocket's height h (in feet) after t seconds is given by the following. h=201t -16t^2 Find all values of for which the rocket's height is 93 feet. Round your answer(s) to the nearest hundredth. (If there is more than one answer, use the "or" button.)

OpenStudy (anonymous):

let h=93 -16t^2+201t=93 -16t^2+201t-93=0 use formular for t t=12.43 seconds or t=0.125second